OpenStudy (anonymous):

solve the equation. -9d-24=-3(d+8)-6d

OpenStudy (uri):

-9d-24=-3(d+8)-6d -9d -24 =-3d -24-6d -9d +3d +6d =-24 +24 -9d +9d=-24 +24 0 =0 They both will be cancelled,equation can't be solved.

OpenStudy (anonymous):

Thank you. Your amazing(: Your are helping me really great. Thanks.